OpenStudy (anonymous):

One winter night the temperature fell 14 degrees between midnight and 5 am. by 9 am the temperature had doubled from what it was at 5 am. by noon, it had risen another 8 degrees to 32 degrees. what was the temperature at midnight?

OpenStudy (anonymous):

TM=Temp. at Midnight Temp 5:00=TM-14 Temp 9:00=2(TM-14) Temp Noon=32=8+2(TM-14) 32=8+2TM-28 32=2TM-20 52=2TM 26=TM